The speed if a car increased from 36km/h to 96km/h in 0.6km.how long did this take and what was the acceleration


1
Expert's answer
2021-03-29T08:57:29-0400

The acceleration of the car by definition is:


"a = \dfrac{v_f-v_i}{t}"

where "v_f = 96km/h = \dfrac{80}{3}m/s" is the final speed of the car, "v_i = 36km\/h = 10m\/s" is the initial speed of the car, and "t" is the time of acceleration.

The distance "d = 0.6km = 600m" travelled under the constant acceleration is given by the kinematic equation:


"d = v_it+\dfrac{at^2}{2}"

Substituting "a", obtain:


"d = v_it + \dfrac{v_ft}{2} - \dfrac{v_it}{2}= \dfrac{(v_f+v_i)t}{2}\\
t = \dfrac{2d}{v_f+v_i} \\
t = \dfrac{2\cdot 600}{80/3+10}\approx 32.7s"

Answer. 32.7 s